OpenStudy (anonymous):

A polynomial function can be written as (x + 1)(x + 4)(x - 7). What are the x-intercepts of the graph of this function? (1, 0), (4, 0), (7, 0) (-1, 0), (-4, 0), (-7, 0) (1, 0), (4, 0), (-7, 0) (-1, 0), (-4, 0), (7, 0)

OpenStudy (anonymous):

@Conqueror i was thinking b but i am not sure

OpenStudy (anonymous):

@bibby @gioskywalker @emmigrace222 @just_one_last_goodbye @AnswerMyQuestions @phi @undeadknight26 @Rainbow_dashie

OpenStudy (anonymous):

@QuantumRanger @wizfan @El_loko_jose @animal_lover36 @stonewoods @HHS6281019 @yungkrystal @leonardo0430

OpenStudy (leonardo0430):

mother of god how many ppl do you know XD

OpenStudy (anonymous):

lol a lot i would guess

OpenStudy (anonymous):

no one knows?

OpenStudy (anonymous):

@Frostbite @justsmile531 @kirbykirby @kevindragonlord @Kittymeowb @kneaux @One098 @Night-Watcher @notsobrainiac @NoDaniel @navvy @Donblue22 @stonewoods @zpupster @zephyr141 @zephyr141 @ZnappyDooZ @ZoMbAsH @ZippySlug @xo.minnie.xox @XVasHappeninZen @xXDNAngelXx

OpenStudy (anonymous):

to find the x-intercepts that mean y = 0 y=(x - 1)(x - 4)(x + 7) 0=(x - 1)(x - 4)(x + 7) so we conclude that on of the multible should be 0 0= x - 1 or 0= x - 4 or 0= x +7 find x values in the three cases.

OpenStudy (anonymous):

does that hlp?

OpenStudy (anonymous):

i think, so your saying its not b? right i thought it might be B

OpenStudy (anonymous):

@xXDNAngelXx

OpenStudy (anonymous):

idk what do you think

OpenStudy (anonymous):

as i said i was thinking B but idk this is why i ask if it is @xXDNAngelXx

OpenStudy (just_one_last_goodbye):

@SouthernRebel101 so u dont get introuble please do not mass tag people u can get suspended for this just advice

OpenStudy (anonymous):

i'm sorry i just want to know the answer and people wont help that much

OpenStudy (anonymous):

Oh wait! Is this what you need? http://www.mathway.com/graph/MzYzNzE

OpenStudy (anonymous):

@SouthernRebel101

OpenStudy (anonymous):

TY SO MUCH

OpenStudy (anonymous):

No Problem :P Mathway.com is a great site

OpenStudy (zephyr141):

XDNangle gave you the answer. it's right there in his/her reply. you have to solve for x in each of those situations.

OpenStudy (zephyr141):

solve for the first one. 0 = x + 1 -1 -1 --------- -1 = x that's the first one. now do the same for the rest of them.

OpenStudy (anonymous):

Oh, okay sorry i thought i had what she was asking. According to all of the websites and videos i watched, i was right but whatever

OpenStudy (anonymous):

there u go @zephyr141 and im a boy

OpenStudy (anonymous):

ty @zephyr141

OpenStudy (anonymous):

(x + 1)(x + 4)(x - 7)=0 iff (x + 1)=0 OR (x + 4)=0 OR (x - 7)=0 implies x=-1 OR x=-4 OR x=7 so answer is (-1, 0), (-4, 0), (7, 0)

OpenStudy (anonymous):

quick question, which grade are you?

OpenStudy (anonymous):

I was right! I just told the wrong answer but showed the right info lol

OpenStudy (anonymous):

@navvy did u see the link i did? Is that the same info u just said?